Compute the numerical value of Gravimetric factor for Magnesium ammonium phosphate hexahydrate (MgNH4PO4.6H2O) precipitate into phosphorous (as tetraphosphorous decoxide, P4O10) .


1
Expert's answer
2021-03-22T06:21:08-0400

Gf = Mass of mg MgNH4PO4.6H2O/ molecular Mass of MgNH4PO4.6H2O + P4O10

Molar mass of MgNH4PO4.6H2O

= 24.3050 + 14.0067 + 30.973762 + 4 (15.994) + 6 (2.01588 + 15.994)

= 245. 4065g/mol

Molecular Mass of molecular Mass of MgNH4PO4.6H2O + P4O10

= 24.3050 + 14.0067 + 30.973762 + 4 (15.994) + 6 (2.01588 + 15.994) + 4(30.973762) + 10(15.994)

= 529.241548

Gf = 245.4065/529.241548

= 0.4637